The length of a rectangular playing field is 5 feet less than twice its width. If the perimeter of the playing field is 230 feet, find the length and width of the field.
Let w= the width of the rectangle and L= length

Answer by checkley71(8403)

L=2W-5
2(2W-5)+2W=230
4W-10+2W=230
6W=230+10
6W=240
W=240/6
W=40 FEET IS THE WIDTH.
L=2*40-5
L=80-5
L=75 FEET IS THE LENGTH.
PROOF
2*75+2*40=230
150+80=230
230=230